OpenStudy (anonymous):

how to find range of: F(X) = 2+3^X-4 :S

OpenStudy (anonymous):

is it 3^(x-4) or (3^x) - 4

OpenStudy (anonymous):

3^(X-4)

OpenStudy (anonymous):

ok thats simple then: First, no exponent will ever give a negative number, so your Range is all positve values Next, no exponent will ever give you a zero either, so its all values greater than zero Last since its f(x) = 2 + 3^(x-4), it must be bigger than 2 since 3^(x-4) must be bigger than zero. Writing it all out \[2 < f(x) < \infty\] or \[(2,\infty)\]